OpenStudy (anonymous):

is this correct >? y=5x+10;(2,14) = y=5x+4

OpenStudy (anonymous):

y=5x+10;(2,14) = y=5x+4

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

y = 5x + 10 .....(2,14)..x = 2 and y = 14 y = mx + b slope(m) = 5 now we sub 14 = 5(2) + b 14 = 10 + b 14 - 10 = b 4 = b equation is : y = 5x + 4 .....yep, you got it